integrate the following integral
∫ (2x + 3)/(x – 1)^2 dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int \frac{2x+3}{(x-1)^2}dx
I = \int \frac{2x-2+5}{(x-1)^2}dx
I = \int [\frac{2}{x-1}+\frac{5}{(x-1)^2}]dx
I = \int \frac{2}{x-1}dx+\int \frac{5}{(x-1)^2}dx
I = 2log(x-1)+\int \frac{5}{(x-1)^2}dx
I = 2log(x-1)-\frac{5}{(x-1)}+c